# Multimodal-Enhanced Objectness Learner for Corner Case Detection in Autonomous Driving

### v1

@@ Abstract
Previous works on object detection have achieved high accuracy in closed-set scenarios, but their performance in open-world scenarios is not satisfactory. One of the challenging open-world problems is corner case detection in autonomous driving. Existing detectors struggle with these cases, relying heavily on visual appearance and exhibiting poor generalization ability. In this paper, we propose a solution by reducing the discrepancy between known and unknown classes and introduce a multimodal-enhanced objectness notion learner. Leveraging both vision-centric and image-text modalities, our semi-supervised learning framework imparts objectness knowledge to the student model, enabling class-aware detection. Our approach, Multimodal-Enhanced Objectness Learner (MENOL) for Corner Case Detection, significantly improves recall for novel classes with lower training costs. By achieving a 76.6% mAR-corner and 79.8% mAR-agnostic on the CODA-val dataset with just 5100 labeled training images, MENOL outperforms the baseline ORE by 71.3% and 60.6%, respectively. The code will be available at https://github.com/tryhiseyyysum/MENOL.
@@ 1 Introduction
In contrast to ZSOD or OVOD, we propose a Multimodal-Enhanced Objectness Learner (MENOL) for Corner Case Detection. MENOL mainly focuses on learning the notion of objectness and improving the recall of novel classes. Many studies have treated the corner case detection problem as an Out-Of-Distribution (OOD) [9] or Anomaly Detection problem [10]. They design very complex rules to detect the corner case objects. However, the t-SNE visualization (Fig.3) shows that the distribution within the novel classes is large, making it difficult to design such general rules. We propose to narrow the gap between known and unknown classes by using geometry cues of depth and normal. To learn the notion of objectness, we design a multimodal-enhanced objectness notion learner, which is a vision-centric and image-text multimodal design. Our MENOL is designed as a semi-supervised framework, using the objectness notion learner as the teacher model to generate class-agnostic pseudo boxes for unlabeled data (the extracted depth and normal images). The pseudo-labeled depth and normal images are then merged with fully-annotated original RGB images and fed into a student model to train the final class-aware open-world object detector. Overall, our proposed MENOL successfully solves the problems mentioned above and significantly improves the recall of novel classes with relatively lower training cost. The idea of learning the notion of objectness and reducing the discrepancy between known and unknown classes is a general way to deal with the problem of corner case detection.
@@
•Vision-centric and image-text multiple modalities are designed to provide additional knowledge and diverse information for the model to learn, enabling it to better understand the world and the notion of objectness.
@@ 3.2 Learning the Notion of Objectness
Since objects in an image are not always in the same scale, the scale variation is a great challenge for autonomous driving object detectors. In our objectness notion learner, the image is first fed into a CNN backbone to extract multi-scale visual features. As shown in Fig.2 Stage II, the extracted features are then fed into the Deformable DETR to obtain vision representation vectors. The attention is calculated at multiple scales by the multi-scale deformable attention module to incorporate better contextual information. To reduce the computational cost, the deformable attention samples a small set of keys around a reference (query) image location to achieve linear complexity with respect to the size of the image feature maps. The pre-trained RoBERTa encodes the text query input and produces a corresponding sequence of hidden vectors. Concatenating the flattened image features and the text embeddings would be likely to destroy the spatial structure of an image. Hence, a better choice is to use a late multimodal fusion mechanism to fuse the image features and the text embeddings. Specifically, the image features are first processed the Deformable DETR architecture to obtain the object query representations. After concatenating with the text embeddings, they are fed into a Transformer architecture to fuse multimodal information. The output head is applied after each transformer self-attention block and the soft token loss and auxiliary losses are calculated to optimize the parameters. Soft token loss is designed for predicting the span of tokens from the original text that refers to each matched object, instead of predicting a categorical class for each detected object. The model is trained to predict a uniform distribution over all token positions that correspond to the object for each predicted box that is matched to a ground truth box using the bi-partite matching. The use of image-text multiple modalities enables the model to learn from large amounts of image-text pairs and get a better understanding of objectness.
@@ 4.1 Datasets and Evaluation Metrics
We conduct training on the combination of CODA [3] and SODA10M [18] dataset. CODA is a novel dataset of object-level corner cases in real world scenes, consisting of approximately 10k carefully selected road driving scenes with high-quality bounding box annotations for 29 representative object categories. SODA10M is a 2D object detection dataset with 6 representative categories. We sample 100 images from CODA dataset and combine them with the SODA10M-labeled training dataset as our training dataset. Then we sample another 488 images from CODA dataset and combine them with the SODA10M-labeled validation dataset as our validation dataset. The model is tested on the CODA-val [3] dataset and BDD100K-val [19] dataset to evaluate its corner case detection and common object detection performance respectively. It simulates the real-world scenario where novel categories will be encountered. We use t-SNE [20] to visualize the distribution of the images in the training dataset. As shown in Fig.3, the pre-trained ResNet-34 is used for feature extraction and the parameters of its fourth layer are extracted for t-SNE dimension reduction. The result shows that the corner cases vary considerably both within and between classes.
@@ 4.3 Implementation Details
The trained objectness notion learner acts as the teacher model to generate pseudo boxes for the extracted depth and normal images in our training dataset (i.e. SODA10M). The pseudo-labeled depth and normal images are merged with fully annotated original RGB images and fed into the student model. The student model is a closed-set object detector based on DINO [17]. It is pre-trained on Objects365 dataset (about 1.7M annotated images) with a Swin Transformer-large backbone. The hyperparameters of the loss function 1 are set as $\lambda_{\text{1}}=2.0$ and $\lambda_{\text{2}}=5.0$. We use AdamW optimizer with initial learning rate of 0.0001 and weight decay of 0.0001 to train the DINO-based student model. The training process accomplishes in 35 epochs using 2 NVIDIA Geforce RTX 3090 GPUs with batch size of 2. Our implementations are based on the tool of MMDetection.
@@ 4.5 Ablation Studies
Finally, we study the effectiveness of our DINO-based [17] student model by comparing it with the Faster R-CNN-based [11] student model. As shown in Fig.6, the DINO-based student model outperforms the Faster R-CNN-based student model, which indicates that the DINO-based student model is more suitable for the task of corner case detection.
@@ 5 Conclusion
In conclusion, our MENOL leverages vision-centric and image-text multimodal learning, incorporating geometry cues to address corner case detection challenges in autonomous driving. By effectively reducing the discrepancy between known and unknown classes, MENOL demonstrates significant improvements in the recall of novel classes, contributing to the robustness of object detection in diverse scenarios. While our approach has shown promising results, future work could explore leveraging richer modalities and more generalized objectness notion learning models to enhance the performance of road detection in autonomous driving.

### v2

@@ Abstract
Previous works on object detection have achieved high accuracy in closed-set scenarios, but their performance in open-world scenarios is not satisfactory. One of the challenging open-world problems is corner case detection in autonomous driving. Existing detectors struggle with these cases, relying heavily on visual appearance and exhibiting poor generalization ability. In this paper, we propose a solution by reducing the discrepancy between known and unknown classes and introduce a multimodal-enhanced objectness notion learner. Leveraging both vision-centric and vision-language multiple modalities, our semi-supervised learning framework imparts objectness knowledge to the student model, enabling class-aware detection. Our approach, Multimodal-Enhanced Objectness Learner (MENOL) for Corner Case Detection, significantly improves recall for novel classes with lower training costs. By achieving a 76.6% mAR-corner and 79.8% mAR-agnostic on the CODA-val dataset with just 5100 labeled training images, MENOL outperforms the baseline ORE by 71.3% and 60.6%, respectively. The code will be available at https://github.com/tryhiseyyysum/MENOL.
@@ 1 Introduction
In contrast to ZSOD or OVOD, we propose a Multimodal-Enhanced Objectness Learner (MENOL) for Corner Case Detection. MENOL mainly focuses on learning the notion of objectness and improving the recall of novel classes. Many studies have treated the corner case detection problem as an Out-Of-Distribution (OOD) [9] or Anomaly Detection problem [10]. They design very complex rules to detect the corner case objects. However, the t-SNE visualization (Fig.3) shows that the distribution difference within the novel classes is large, making it difficult to design such general rules. We propose to narrow the gap between known and unknown classes by using geometry cues of depth and normal. To learn the notion of objectness, we design a multimodal-enhanced objectness notion learner, which is a vision-centric and vision-language multimodal design. Our MENOL is designed as a semi-supervised framework, the objectness notion learner acts as the teacher model to generate class-agnostic pseudo boxes for unlabeled data (the extracted depth and normal images). The pseudo-labeled depth and normal images are then merged with fully-annotated original RGB images and fed into a student model to train the final class-aware open-world object detector. Overall, our proposed MENOL successfully solves the problems mentioned above and significantly improves the recall of novel classes with relatively lower training cost. The idea of learning the notion of objectness and reducing the discrepancy between known and unknown classes is a general way to deal with the problem of corner case detection.
@@
•Vision-centric and vision-language multiple modalities are designed to provide additional knowledge and diverse information for the model to learn, enabling it to better understand the world and the notion of objectness.
@@ 3.2 Learning the Notion of Objectness
Since objects in an image are not always in the same scale, the scale variation is a great challenge for autonomous driving object detectors. In our objectness notion learner, the image is first fed into a CNN backbone to extract multi-scale visual features. As shown in Fig.2 Stage II, the extracted features are then fed into the Deformable DETR to obtain vision representation vectors. The attention is calculated at multiple scales by the multi-scale deformable attention module to incorporate better contextual information. To reduce the computational cost, the deformable attention samples a small set of keys around a reference (query) image location to achieve linear complexity with respect to the size of the image feature maps. The pre-trained RoBERTa encodes the text query input and produces a corresponding sequence of hidden vectors. Concatenating the flattened image features and the text embeddings would be likely to destroy the spatial structure of an image. Hence, a better choice is to use a late multimodal fusion mechanism to fuse the image features and the text embeddings. Specifically, the image features are first processed the Deformable DETR architecture to obtain the object query representations. After concatenating with the text embeddings, they are fed into a Transformer architecture to fuse multimodal information. The output head is applied after each transformer self-attention block and the soft token loss and auxiliary losses are calculated to optimize the parameters. Soft token loss is designed for predicting the span of tokens from the original text that refers to each matched object, instead of predicting a categorical class for each detected object. The model is trained to predict a uniform distribution over all token positions that correspond to the object for each predicted box that is matched to a ground truth box using the bi-partite matching. The use of vision-language multiple modalities enables the model to learn from large amounts of image-text pairs and get a better understanding of objectness.
@@ 4.1 Datasets and Evaluation Metrics
We conduct training on the combination of CODA [3] and SODA10M [18] dataset. CODA is a novel dataset of object-level corner cases in real world scenes, consisting of approximately 10k carefully selected road driving scenes with high-quality bounding box annotations for 29 representative object categories. SODA10M is a 2D object detection dataset with 6 representative categories. We sample 100 images from CODA dataset and combine them with the SODA10M-labeled training dataset as our training dataset. Then we sample another 488 images from CODA dataset and combine them with the SODA10M-labeled validation dataset as our validation dataset. The model is tested on the CODA-val [3] dataset (not overlap with the aforementioned datasets) and BDD100K-val [19] dataset to evaluate its corner case detection and common object detection performance respectively. It simulates the real-world scenario where novel categories will be encountered. We use t-SNE [20] to visualize the distribution of the images in the training dataset. As shown in Fig.3, the pre-trained ResNet-34 is used for feature extraction and the parameters of its fourth layer are extracted for t-SNE dimension reduction. The result shows that the corner cases vary considerably both within and between classes.
@@ 4.3 Implementation Details
The trained objectness notion learner acts as the teacher model to generate pseudo boxes for the extracted depth and normal images in our training dataset (i.e. SODA10M). The pseudo-labeled depth and normal images are merged with fully annotated original RGB images and fed into the student model. The student model is a closed-set object detector based on DINO [17]. It is pre-trained on Objects365 dataset (about 1.7M annotated images) with a Swin Transformer-large backbone. The hyperparameters of the loss function 1 are set as $\lambda_{\text{1}}=2.0$ and $\lambda_{\text{2}}=5.0$. We use AdamW optimizer with initial learning rate of 0.0001 and weight decay of 0.0001 to train the DINO-based student model. The training process accomplishes in 35 epochs using 2 NVIDIA Geforce RTX 3090 GPUs with batch size of 2.
@@ 4.5 Ablation Studies
Finally, we study the effectiveness of our DINO-based [17] student model by comparing it with Faster R-CNN-based [11] student model. As shown in Fig.6, DINO-based student model outperforms Faster R-CNN-based student model, which indicates that DINO-based student model is better.
@@ 5 Conclusion
In conclusion, our MENOL leverages vision-centric and vision-language multimodal learning, incorporating geometry cues to address corner case detection challenges in autonomous driving. By effectively reducing the discrepancy between known and unknown classes, MENOL demonstrates significant improvements in the recall of novel classes. While our approach has shown promising results, future work could explore richer modalities and more generalized objectness notion learning models to enhance the performance of road detection in autonomous driving.